# Evaluating link prediction: New perspectives and recommendations

### v1

@@ Abstract
Link prediction (LP) is an important problem in network science and machine learning research. The state-of-the-art LP methods are usually evaluated in a uniform setup, ignoring several factors associated with the data and application specific needs. We identify a number of such factors, such as, network-type, problem-type, geodesic distance between the end nodes and its distribution over the classes, nature and applicability of LP methods, class imbalance and its impact on early retrieval, evaluation metric, etc., and present an experimental setup which allows us to evaluate LP methods in a rigorous and controlled manner. We perform extensive experiments with a variety of LP methods over real network datasets in this controlled setup, and gather valuable insights on the interactions of these factors with the performance of LP through an array of carefully designed hypotheses. Following the insights, we provide recommendations to be followed as best practice for evaluating LP methods.
@@
This paper is organized as follows. In Section 2, we summarize the related works. We provide the details of the LP methods which we use in our experiments in Section 3, grouped by their categories. Section 4 provides the details of the datasets on which we apply the LP methods. We describe the proposed experimental setup in Section 5. In Section 6, we present the results and analyze our findings, and provide our recommendations. We conclude our work and indicate the future directions in Section 7.
@@ 2 Related Works
Nowell et al. [1] proposed a list of local and global similarity based LP methods. Similarity based methods calculate a score between two nodes to estimate their chance to form a link. Local similarity based methods are neighborhood based, and apply to two-hop away node pairs. The global similarity based methods are mostly path based and apply to any node pair in the network. In this paper, we consider most of the methods proposed in Nowell et al. in our experiments along with another popular local method: Resource Allocation [17]. There are a few other similarity based methods in literature, an exhaustive list of which can be found in Kumar et al. [2]. We do not include another type of similarity based method, namely quasi-local [2], in our study. Modern machine learning based LP methods are either supervised [18, 27, 28, 29] or unsupervised [19, 20, 21, 22]. The supervised methods train a neural network or a graph neural network in an end-to-end manner with the test node pairs. The unsupervised methods first train a neural network or a graph neural network to learn low dimension embeddings of the nodes in an unsupervised manner (known as contrastive learning), and then produce link features which are used to build a classifier to predict links. In this paper, we consider the unsupervised methods in our experiments. The main reason to exclude the end-to-end methods is the difficulty in building the experimental set-up for those in the longitudinal setup for the future LP, which incurs a longitudinal bias [3, 35] on the prediction performance, as well as makes the methods incomparable with their missing LP counterpart. Another group of unsupervised LP approaches are matrix factorization based [23, 36], which we do not include in our experiments. Note that, the purpose of this work is not to present an exhaustive performance comparison of the existing methods, but to devise a robust evaluation strategy for the prediction methods considering certain data, evaluation metric, experimental setup, application and methodology specific factors. So, excluding a few prediction methods does not hurt the merit of this work.
@@
The start-of-the-art LP methods briefed above mainly apply to a simple, undirected and homogeneous networks. There exists works which exploit several network attributes to predict links. Few examples are: temporal link prediction [7, 25] which exploits temporal dynamics of link maintenance in dynamic networks like in social networks; link prediction in bipartite networks [37, 38] which applies on bipartite graphs like user-product networks or term-document networks; link prediction in heterogeneous networks [7, 39, 40] which exploits the heterogeneity of nodes and links in networks like bibliographic networks; link prediction in directed networks [41, 42] which exploits the direction information in directed networks like email networks, etc. These specialized methods are not necessarily disjoint, and often use information covering multiple groups. In this paper, we consider the start-of-the-art LP methods only in our analysis, and reserve the specialized methods for future study.
@@ 3.2.5 Naming convention of individual link prediction methods:
4 Datasets
We conducted our experiments on fifteen real network datasets of diverse nature, such as, social network, communication network, collaboration network, biological network, power grid, etc. A few of these networks like friendship networks, collaboration networks are undirected in nature, whereas networks like communication networks are directed. The social, communication and collaboration networks are inherently dynamic or temporal in nature. Among the networks in our collection which are dynamic, few had timestamps available with the interactions or the moment of link formation. In total, nine among all the networks had timestamps. For these networks we created ground truth labels two ways: (a) considering the timestamps which lets us evaluate future LP, and (b) removing a portion of the links in the networks which lets us evaluate the missing LP. So, we created two LP datasets for each of these networks. For the networks with no timestamp, we could only evaluate missing LP. In total, we conducted our experiments and analysis on datasets, among which 15 were used for missing LP and the rest were used for future LP. We summarize the datasets in Table 2. We briefly describe the datasets as follows.
4.1 Networks with timestamps:
1.
2.
3.
4.
5.
6.
7.
8.
9.
4.2 Networks without timestamps:
1.
2.
3.
4.
5.
6.
@@ 4.2 Networks without timestamps:
| CollegeMsg | 1899 | 20296 | Directed | 194 days | clg-msg-t | 2142 | 1068(49.80%) | 983(45.89%) |
@@
| Email-Eu | 986 | 332334 | Directed | 803 days | email-t | 3977 | 3774(94.89%) | 200(5.03%) |
@@
| Mathoverflow | 24818 | 506550 | Directed | 2350 days | math-flow-t | 24225 | 18918(78.09%) | 4889(20.18%) |
@@
| Ast-Phys | 9756 | 86697 | Undirected | 3316 days | ast-phys-t | 32186 | 27035(83.99%) | 5151(16%) |
@@
| LastFm | 7624 | 27806 | Undirected | 1345 days | lst-fm-t | 29867 | 29862(99.38%) | 185(0.61%) |
@@
| IAcontact | 113 | 2498 | Directed | 1 day | iacontact-t | 533 | 533(100%) | – |
@@
| FB-forum | 899 | 7089 | Directed | 164 days | forum-t | 1638 | 717(43.58%) | 862(52.81%) |
@@
| Topology | 34761 | 171403 | Undirected | 23 days | topo-t | 21100 | 17331(82.13%) | 3159(14.97%) |
@@
| Mooc | 7047 | 411749 | Directed | 29 days | act-mooc-t | 38448 | 38381(99.82%) | 67(0.01%) |
@@
| ArXiv | 23133 | 93133 | Undirected | arxiv | 886 | 867(97.85%) | 6(0.14%) | |
| Power-grid | 4941 | 6594 | Undirected | pow-grid | 183 | 117(63.93%) | 65(35.51%) | |
| Tech-routers | 2113 | 6632 | Undirected | routers | 590 | 72(12.20%) | 45(7.62%) | |
| Bio-yeast | 6008 | 156945 | Undirected | bio-yeast | 1572 | 1472(93.63%) | 100(6.36%) | |
| 4039 | 88234 | Undirected | fb | 8813 | 8796(99.80%) | 17(0.19%) | ||
| BlogCatalog | 10312 | 333983 | Undirected | blog-cat | 33350 | 33112(99.28%) | 238(0.71%) |
@@ 5.1 Graph and ground truth generation
In this section, we detail the procedure of how we prepare the graph and the ground truth positive and negative node pairs from the datasets presented in Section 4, for future and missing LP tasks. Unlike traditional approaches, we carefully segregate the test examples based on the distance of the test node pairs in the graph, and maintain the distance bias inherent with the positive examples’ distribution, while sampling the negative examples. Moreover, it takes care of the evaluation in a class imbalance environment and its interactions with the factors stated above.
The future LP datasets were prepared for the networks where the interaction or link formation timestamps were available. The interactions can be directed, and there could be multiple interactions between a node pair. As we aim to prepare a simple and undirected graph and the true future links from a network, we ignored the directions of the interactions and the duplicate interactions. The detailed procedure for the dataset preparation is described as follows. We first sort the interactions or links (for the cases where only the link formation timestamps were available) by their timestamp values, and then choose of the earlier timestamped interactions or links for constructing the graph, and the remaining were used for generating positive examples. From the interactions which are in that , we constructed an unweighted and undirected graph by first ignoring the directions and then removing duplicate interactions and links, if any. If the graph is having multiple connected components, we choose the largest connected component as the final graph for our experiment. From the remaining interactions, we ignored the directions and removed the duplicates to generate the potential positive examples. From this potential set, we first ignored the node pairs which involve at least one node which is not included in the graph, and then removed the links which are present in the graph as an edge to get the final set of positive examples.
The missing LP was applied to all the networks presented in Section 4. We preprocessed the networks having timestamp information by ignoring the timestamps and their directions, and then removing the duplicates. Similarly, we ignored the directions and removed the duplicates, if applicable, for the networks which did not have timestamp information. After doing this, we are left with a set of node pairs which represents links. Then we randomly remove of the links from the set. We construct the graph with the remaining links and choose the largest connected component as the final graph. We consider the removed links as the potential positive examples, and generate the positive examples following the same procedure as the future LP datasets.
@@ 5.3 Evaluation measures
The area under the receiver operating curve (AUROC) [31] has historically been the primary performance metric used to assess the performance of LP methods [2, 24]. The receiver operating curve (ROC) plots the true positive rate (TPR) against the false positive rate (FPR) at every threshold value. AUROC gives the probability of a positive example assigned a higher score than that of a negative example. However, AUROC is not a suitable performance measure in a class imbalance scenario like LP problem, which is inherently skewed towards the negative class. AUROC favors accurate classification of positive examples, at the cost of significantly misclassifying negative examples. Hence, AUROC can overestimate the performance of LP. An alternative to AUROC is area under the precision-recall curve (AUPR) [32, 3]. Precision-recall (PR) curve plots precision at every threshold value vs. its recall, i.e., TPR. As precision is sensitive to class imbalance unlike FPR, it can provide better evaluation of LP [32]. Despite AUPR’s suitability, it is less popular in LP literature, may be due to the variability of PR curves in different skew levels, which makes it difficult to compare the LP methods when multiple skewness is considered. To deal with class imbalance, another skew sensitive evaluation measure, is sometimes used to evaluate LP as a single threshold measure, where the positive example count is considered as a typical value of [65, 2].
@@ 6 Results and Discussion
We applied the 58 LP methods presented in Section 3 on the 24 datasets defined in Section 4 in a controlled experimental setup proposed in Section 5. We evaluated their performance based on the four evaluation metrics: AUROC, AUPR, and on the test sets. As described in Section 5.1, we had different test sets depending on the shortest path distance between test node pairs, and each of these had two versions: balanced and imbalanced. For the skew sensitive metrics: AUPR, and , we consider evaluating both for balanced and imbalanced test set. We performed a series of statistical tests to test an array of carefully designed hypotheses in this controlled setup to interrogate how the LP methods perform in different situations depending on the shortest path distance between test node pairs, prediction type, network type, and class imbalance. We also investigate how the class imbalance affect the LP methods in early retrieval phase using proper statistical tests and PRC plots. We present the detailed results along with the insights acquired from them as follows.
@@
In this section, we investigate whether the prediction performance of the LP methods differ on a particular network for the two cases: future LP and missing LP. Given a network, we apply paired t-test to compare the absolute values of the prediction performances of the prediction methods given a particular evaluation metric for the two cases. To find the level of the ordinal association between the two cases, we apply Kendall?s Tau Rank Correlation Coefficient (). We consider AUROC, and both balanced and imbalanced versions of AUPR and as the evaluation measure. We perform the experiment on the nine networks which had timestamps in distance controlled test set. Table 3 and Table 4 summarize the results of the two tests respectively.
Almost all paired t-test p-values are significant across distances and evaluation measures, and for almost all networks, the LP methods perform better for future LP than missing LP. The only exception is the two-hop case of LastFm network, where either the missing LP wins or there is no significant difference. When the test node pairs are generated randomly in a longitudinal network, the temporal relationship among the links are lost. The outperformance of future LP might be due to the adverse affect of this information loss in their missing LP counterparts. This also indicates that when an application demands future LP, but the ground truth links are generated in traditional way, i.e., removing a portion of links from the network randomly, the prediction performance is significantly underestimated, and vice versa. Table 4 shows that except for the Topology and Mooc network, all rank correlation values are extremely low, in few cases even negative. This indicates that, for a future LP, if the ground truth links are generated in traditional way, the prediction performance rankings of various LP methods vary significantly than if the ground truth links were generated with true future links. Our results indicate the necessity to evaluate a LP method in both setups and report its performance in both types of prediction tasks, in order to understand its effectiveness.
@@
| NETWORK | AUROC | AUPR | Pr@P | |||
| balanced | imbalanced | balanced | imbalanced | |||
| two-hop | CollegeMsg | 0.184 | 0.184 | 0.219 | 0.156 | 0.203 |
| Email-Eu | 0.272 | 0.319 | 0.472 | 0.294 | 0.526 | |
| Mathoverflow | 0.339 | 0.386 | 0.414 | 0.326 | 0.488 | |
| Ast-phys | 0.487 | 0.433 | 0.377 | 0.525 | 0.393 | |
| lastFm | 0.477 | 0.465 | 0.381 | 0.462 | 0.335 | |
| IAcontact | 0.027 | 0.080 | 0.038 | 0.039 | 0.075 | |
| FB-forum | 0.321 | 0.347 | 0.438 | 0.311 | 0.271 | |
| Topology | 0.785 | 0.783 | 0.771 | 0.723 | 0.742 | |
| Mooc | 0.744 | 0.743 | 0.698 | 0.678 | 0.741 | |
| three-hop | CollegeMsg | 0.009 | 0.118 | 0.158 | 0.022 | 0.230 |
| Email-Eu | -0.030 | 0.085 | 0.167 | -0.144 | 0.181 | |
| Mathoverflow | -0.020 | -0.071 | 0.018 | 0.139 | 0.060 | |
| Ast-phys | 0.256 | 0.464 | 0.578 | 0.422 | 0.628 | |
| lastFm | — | — | — | — | — | |
| IAcontact | — | — | — | — | — | |
| FB-forum | 0.189 | 0.208 | 0.283 | 0.203 | 0.334 | |
| Topology | 0.749 | 0.747 | 0.738 | 0.649 | 0.773 | |
| Mooc | — | — | — | — | — | |

### v2

@@ Abstract
Link prediction (LP) is an important problem in network science and machine learning research. The state-of-the-art LP methods are usually evaluated in a uniform setup, ignoring several factors associated with the data and application specific needs. We identify a number of such factors, such as, network-type, problem-type, geodesic distance between the end nodes and its distribution over the classes, nature and applicability of LP methods, class imbalance and its impact on early retrieval, evaluation metric, etc., and present an experimental setup which allows us to evaluate LP methods in a rigorous and controlled manner. We perform extensive experiments with 58 LP methods of diverse categories over 24 real network datasets in this controlled setup, and gather valuable insights on the interactions of these factors with the performance of LP through an array of carefully designed hypothesis tests. Following the insights, we provide recommendations on the experimental setup and the evaluation measures to be employed considering application scenarios, to be followed as best practice for evaluating LP methods.
@@
This paper is organized as follows. In Section 2, we summarize the related works. We provide the details of the LP methods which we use in our experiments in Section 3, grouped by their categories. Section 4 provides the details of the datasets on which we apply the LP methods along with the proposed experimental setup. In Section 5, we present the results and analyze our findings, and provide our recommendations. We conclude our work and indicate the limitations and future directions in Section 6.
@@ 2 Related Works
Nowell et al. [1] proposed a list of local and global similarity based LP methods. Similarity based methods calculate a score between two nodes to estimate their chance to form a link. Local similarity based methods are neighborhood based, and apply to two-hop away node pairs. The global similarity based methods are mostly path based and apply to any node pair in the network. In this paper, we consider most of the methods proposed in Nowell et al. in our experiments along with another popular local method: Resource Allocation [17]. There are a few other similarity based methods in literature, an exhaustive list of which can be found in Kumar et al. [2]. We do not include another type of similarity based method, namely quasi-local [2], in our study. Modern machine learning based LP methods are either supervised [18, 27, 28, 29] or unsupervised [19, 20, 21, 22]. The supervised methods train a neural network or a graph neural network in an end-to-end manner with the test node pairs. The unsupervised methods first train a neural network or a graph neural network to learn low dimension embeddings of the nodes in an unsupervised manner (known as contrastive learning), and then produce link features which are used to build a classifier to predict links. In this paper, we consider the unsupervised methods in our experiments. The main reason to exclude the end-to-end methods is the difficulty in building the experimental set-up for those in a longitudinal setup, for the future LP, which incurs a longitudinal bias [3, 35] on the prediction performance, as well as makes the methods incomparable with their missing LP counterpart. Another group of unsupervised LP approaches are matrix factorization based [23, 36], which we do not include in our experiments. Note that, the purpose of this work is not to present an exhaustive performance comparison of the existing methods, but to devise a robust evaluation strategy for the prediction methods considering certain data, evaluation metric, experimental setup, application and methodology specific factors. So, excluding a few prediction methods does not hurt the merit of this work.
@@
The state-of-the-art LP methods briefed above mainly apply to a simple, undirected and homogeneous networks. There exist works which exploit several network attributes to predict links. Few examples are: temporal link prediction [7, 25] which exploits temporal dynamics of link maintenance in dynamic networks like in social networks; link prediction in bipartite networks [37, 38] which applies on bipartite graphs like user-product networks or term-document networks; link prediction in heterogeneous networks [7, 39, 40] which exploits the heterogeneity of nodes and links in networks like bibliographic networks; link prediction in directed networks [41, 42] which exploits the direction information in directed networks like email networks, etc. These specialized methods are not necessarily disjoint, and often use information covering multiple groups. In this paper, we consider the state-of-the-art LP methods only in our analysis, and reserve the specialized methods for future study.
@@ 3.2.5 Naming convention of individual link prediction methods:
| CollegeMsg [57] | 1899 | 20296 | Directed | 194 days | clg-msg-t | 2142 | 1068(49.80%) | 983(45.89%) |
@@
| Email-Eu [58] | 986 | 332334 | Directed | 803 days | email-t | 3977 | 3774(94.89%) | 200(5.03%) |
@@
| Mathoverflow [58] | 24818 | 506550 | Directed | 2350 days | math-flow-t | 24225 | 18918(78.09%) | 4889(20.18%) |
@@
| Ast-Phys [59] | 9756 | 86697 | Undirected | 3316 days | ast-phys-t | 32186 | 27035(83.99%) | 5151(16%) |
@@
| LastFm [60] | 7624 | 27806 | Undirected | 1345 days | lst-fm-t | 29867 | 29862(99.38%) | 185(0.61%) |
@@
| IAcontact [61] | 113 | 2498 | Directed | 1 day | iacontact-t | 533 | 533(100%) | – |
@@
| FB-forum [61] | 899 | 7089 | Directed | 164 days | forum-t | 1638 | 717(43.58%) | 862(52.81%) |
@@
| Topology [62, 63] | 34761 | 171403 | Undirected | 23 days | topo-t | 21100 | 17331(82.13%) | 3159(14.97%) |
@@
| Mooc [60] | 7047 | 411749 | Directed | 29 days | act-mooc-t | 38448 | 38381(99.82%) | 67(0.01%) |
@@
| ArXiv [59] | 23133 | 93133 | Undirected | arxiv | 886 | 867(97.85%) | 6(0.14%) | |
| Power-grid [61] | 4941 | 6594 | Undirected | pow-grid | 183 | 117(63.93%) | 65(35.51%) | |
| Tech-routers [61] | 2113 | 6632 | Undirected | routers | 590 | 72(12.20%) | 45(7.62%) | |
| Bio-yeast [61] | 6008 | 156945 | Undirected | bio-yeast | 1572 | 1472(93.63%) | 100(6.36%) | |
| Facebook [64] | 4039 | 88234 | Undirected | fb | 8813 | 8796(99.80%) | 17(0.19%) | |
| BlogCatalog [61] | 10312 | 333983 | Undirected | blog-cat | 33350 | 33112(99.28%) | 238(0.71%) |
@@ 4 Experimental Setup
4.1 Network datasets
We conducted our experiments on fifteen real network datasets of diverse nature, such as social network, communication network, collaboration network, biological network, power grid, etc. A few of these networks like friendship networks, collaboration networks are undirected in nature, whereas networks like communication networks are directed. The social, communication and collaboration networks are inherently dynamic or temporal in nature. Among the networks in our collection which are dynamic, few had timestamps available with the interactions or the moment of link formation. In total, nine among all the networks had timestamps. For these networks we created ground truth labels in two ways: (a) considering the timestamps which lets us evaluate future LP, and (b) removing a portion of the links in the networks which lets us evaluate the missing LP. So, we created two LP datasets for each of these networks. For the networks with no timestamp, we could only evaluate missing LP. In total, we conducted our experiments and analysis on datasets, among which 15 were used for missing LP and the rest were used for future LP. We summarize the datasets in Table 2. Note that, the column titled “Network” gives the network names, where the “Dataset” column gives the names of the datasets for link prediction after generating the graph and ground truth labels. The first nine networks had timestamps, so each produces two datasets. The name of the datasets which consider timestamps, are suffixed with -t. For an example, the “CollegeMsg” network produces two datasets; the timestamped one is named as clg-msg-t, and the non-timestamped one is named as clg-msg. We present a brief description of the networks at Appendix A.
@@ 4.2 Graph and ground truth generation
In this section, we detail the procedure of how we prepare the graph and the ground truth positive and negative node pairs from the network datasets presented in Section 4.1, for future and missing LP tasks. Unlike traditional approaches, we carefully segregate the test examples based on the distance of the test node pairs in the graph, and maintain the distance bias inherent with the positive examples’ distribution, while sampling the negative examples. Moreover, it takes care of the evaluation in a class imbalance environment and its interactions with the factors stated above.
The future LP datasets were prepared for the networks where the interaction or link formation timestamps were available. The interactions can be directed, and there could be multiple interactions between a node pair. As we aim to prepare a simple and undirected graph and the true future links from a network, we ignored the directions of the interactions and the duplicate interactions. The detailed procedure for the dataset preparation is described as follows. We first sort the interactions or links (for the cases where only the link formation timestamps were available) by their timestamp values, and then choose of the earlier timestamped interactions or links for constructing the graph, and the remaining were used for generating positive examples. From the interactions which are in that , we constructed an unweighted and undirected graph by first ignoring the directions and then removing duplicate interactions and links, if any. If the graph is having multiple connected components, we choose the largest connected component as the final graph for our experiment. From the remaining interactions, we ignored the directions and removed the duplicates to generate the potential positive examples. From this potential set, we first ignored the node pairs which involve at least one node which is not included in the graph, and then removed the links which are present in the graph as an edge to get the final set of positive examples. A pseudocode for the graph and positive examples generation for future LP is given in Algorithm 1.
training_interactions first 75%
future_interactions last 25%
The missing LP was applied to all the networks presented in Section 4.1. We preprocessed the networks having timestamp information by ignoring the timestamps and their directions, and then removing the duplicates. Similarly, we ignored the directions and removed the duplicates, if applicable, for the networks which did not have timestamp information. After doing this, we are left with a set of node pairs which represents links. Then we randomly remove of the links from the set. We construct the graph with the remaining links and choose the largest connected component as the final graph. We consider the removed links as the potential positive examples, and generate the positive examples following the same procedure as the future LP datasets. A pseudocode for the graph and positive examples generation for missing LP is given in Algorithm 2.
@@ 4.4 Evaluation measures
The area under the receiver operating curve (AUROC) [31] has historically been the primary performance metric used to assess the performance of LP methods [2, 24]. The receiver operating curve (ROC) plots the true positive rate (TPR) against the false positive rate (FPR) at every threshold value. AUROC gives the probability of a positive example assigned a higher score than that of a negative example. However, AUROC is not a suitable performance measure in a class imbalance scenario like LP problem, which is inherently skewed towards the negative class. AUROC favors accurate classification of positive examples, at the cost of significantly misclassifying negative examples. Hence, AUROC can overestimate the performance of LP. An alternative to AUROC is area under the precision-recall curve (AUPR) [32, 3]. Precision-recall (PR) curve plots precision at every threshold value vs. its recall, i.e., TPR. As precision is sensitive to class imbalance unlike FPR, it can provide better evaluation of LP [32]. Despite AUPR’s suitability, it is less popular in LP literature, may be due to the variability of PR curves in different skew levels, which makes it difficult to compare the LP methods when multiple skewness is considered. Measures like , , F-score, etc. are less popular in link prediction literature, mainly because, unlike AUROC and AUPR, they measure performance for a single threshold value, which prohibits us to understand the method’s behavior throughout the whole spectrum of thresholds [3, 2, 24]. , another skew sensitive evaluation measure, is sometimes used to evaluate LP as a single threshold measure, where the positive example count is considered as a typical value of [65, 2].
@@ 5 Results and Discussion
We applied the 58 LP methods presented in Section 3 on the 24 datasets defined in Section 4.1 in a controlled experimental setup proposed in Section 4. We evaluated their performance based on the evaluation metrics presented in Section 4.4. We performed a series of statistical tests to test an array of carefully designed hypotheses in this controlled setup to interrogate how the LP methods perform in different situations depending on the shortest path distance between test node pairs, prediction type, network type, and class imbalance. We also investigate how the class imbalance affect the LP methods in early retrieval phase using proper statistical tests and PRC plots. We present the detailed results along with the insights acquired from them, and our recommendations as follows.
@@
In this section, we investigate whether the prediction performance of the LP methods differ on a particular network for the two cases: future LP and missing LP. Given a network, we apply paired t-test to compare the absolute values of the prediction performances of the prediction methods given a particular evaluation metric for the two cases. To find the level of the ordinal association between the two cases, we apply Kendall’s Tau Rank Correlation Coefficient (). We consider AUROC, and both balanced and imbalanced versions of AUPR and as the evaluation measure. We perform the experiment on the nine networks which had timestamps in distance controlled test set.
Table 3 summarizes the paired t-test results. Almost all paired t-test p-values are significant across distances and evaluation measures, and for almost all networks, the LP methods perform better for future LP than missing LP. The only exception is the two-hop case of LastFm network, where either the missing LP wins or there is no significant difference. When the test node pairs are generated randomly in a longitudinal network, the temporal relationship among the links are lost. The outperformance of future LP might be due to the adverse affect of this information loss in their missing LP counterparts. This also indicates that when an application demands future LP, but the ground truth links are generated in traditional way, i.e., removing a portion of links from the network randomly, the prediction performance is significantly underestimated, and vice versa.
Kendall’s Tau results reveal that except for the Topology and Mooc network, all rank correlation values are extremely low, in few cases even negative. This indicates that, for a future LP, if the ground truth links are generated in traditional way, the prediction performance rankings of various LP methods vary significantly than if the ground truth links were generated with true future links. All correlation values can be found in Table 16 in Appendix C.
Our results indicate the necessity to evaluate a LP method in both setups and report its performance in both types of prediction tasks, in order to understand its effectiveness.